The perimeter of a rectangle is 160 ft. One fourth the length is the same as twice the width. Find the dimensions of the rectangle.
Answer by checkley77(12844)

L/4=2W OR L=8W
2L+2W=160
2*8W+2W=160
16W+2W=160
18W=160
W=160/18
W=80/9 ANS.
L=8*80/9
L=640/9 ANS.
PROOF:
2*640/9+2*80/9=160
1280/9+160/9=160
1440/9=160
160=160